Two capacitors each of capacity C are connected in parallel. The equivalent capacity is 

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student 
Two capacitor in parallel 
Ceq = C + C = 2C 
Hope tis will help …..............................
Good Luck 
Cheers 
Arun
4 years ago
total or equivalent capacitanceCT of an electrical circuit containing two or more Capacitors in Parallel is the sum of the all the individual capacitance's added together as the effective area of the plates is increased.
 
Hence equivalent capacitance = c + c = 2c